OpenStudy (anonymous):

Please help! Simplify: 4x-32y divided by 4

OpenStudy (abb0t):

factor out the 4 from the numerator. You get \(\huge \frac{4(x+8y)}{4}\)

OpenStudy (abb0t):

Remember: \(\large \frac{4}{4}=1\)

OpenStudy (anonymous):

Shouldn't that be \[\frac{ 4(x-8y) }{ 4 }\].

OpenStudy (xanthe):

yes!

OpenStudy (xanthe):

lol

OpenStudy (abb0t):

It's the same thing!

OpenStudy (xanthe):

But nt e signXD